### v2

@@ 3.2 Spectral Diagnostics
A known problem with deterministic ML weather models is that their forecasts are "blurry" (Kochkov et al., 2023). Compared to ERA5, they have reduced power at small wavelengths. Deterministic ML models are often trained using the MSE loss function, which strongly penalizes sharp forecasts in the wrong place. This is referred to as the double penalty problem (Mittermaier, 2014), in which an ensemble is penalized once for predicting a storm in the wrong place and another time for missing the correct location of the storm. To avoid the double penalty from the mean squared error, ML models may learn to predict smooth, blurred solutions. Existing work has noted that these smooth ML predictions appear closer to an ensemble mean (Agrawal et al., 2023; Brenowitz et al., 2024), rather than an individual ensemble member.
@@
However, a key feature of the ensemble mean is not just that the forecasts are blurry. The spectra must increasingly blur with lead time (Bonavita, 2023). As the trajectories diverge during the forecast rollout, the ensemble mean should become blurrier, and its spectra should increasingly degrade at small wavelengths. In particular, on synoptic time scales (around 3-5 days), when error growth becomes nonlinear, the IFS ensemble mean displays a sharp decline in power around 1000 kilometer wavelengths (Bonavita, 2023). This introduces two important criteria for an ML-based weather forecast: whether the forecasts are blurry, and whether they get increasingly blurry with lead time.
On the first criterion, SFNO-BVMC ensemble members do contain some blurriness, like many other deterministic ML models. On the second criterion, crucially, their spectra remain constant through the 360-hour rollout (Figure 10 and Figure 11). This contrasts with GraphCast and AIFS; those deterministic ML models do increasingly blur with lead time (Kochkov et al., 2023; Lang et al., 2024). Brenowitz et al. (2024) and Lang et al. (2024) hypothesize that autoregressive fine-tuning could be responsible for this behavior. In autoregressive fine-tuning, the ML model weights are optimized over multiple timesteps. Normally, during training, the model weights are optimized to minimize the MSE of just 1 timestep of the forecast, but during the autoregressive fine-tuning phase, the weights are optimized based on the predictions and ground truth after multiple time steps. The goal of this method is to improve the performance during rollout. Autoregressive fine-tuning may effectively increase the time step of the model, making it more similar to an ensemble mean (Lang et al., 2024) and contributing to increased blurring during the rollout. To minimize this spectral degradation, we do not conduct any autoregressive fine-tuning. The SFNO trained here is only trained to predict six hours ahead, and its autoregressive performance is not optimized. With this design decision, we reduce the training computational requirements, as autoregressive fine-tuning is intensive in GPU memory and computation time. We use the savings from this choice to train an SFNO with a small scale factor and large embedding dimension. These design choices allow our configuration of SFNO to hold as much high-resolution information in its internal representation as possible.
@@
While the control and perturbed spectra remain constant through the rollout, the SFNO-BVMC ensemble mean does increasingly blur with lead time. Figure 12 shows that the ensemble means of SFNO-BVMC and IFS ENS similarly degrade in power after 24 hours, 120 hours, and 240 hours. For Z500, there is a notable decline in power between lead times of 24 hours and 120 hours. This sharp decline is due to the nonlinear error growth that characterizes forecasts at lead times of 3–5 days. On synoptic scales ($\sim$1000 km in space and 3–5 days in time), SFNO-BVMC’s ensemble mean has a similar decline in power as IFS ENS. This increases our trust that the ensemble members trajectories realistically diverge, and the ensemble is faithfully representing synoptic error growth. Bonavita (2023) originally posed this test comparing the spectra of a deterministic PanGu ML model and the IFS ensemble mean. Despite the blurring in PanGu, they show that a control run of PanGu does not successfully mimic the IFS ensemble mean spectrum.
Through these spectral diagnostics, we validate that the individual members’ predictions do not collapse into the ensemble mean. This is a crucial test of the physical fidelity of SFNO-BVMC. Because each SFNO-BVMC ensemble member’s spectrum is constant through the rollout, the ensemble members maintain their ability to predict extreme weather. If their spectra degraded with lead time, then the forecasts would become too blurry to predict localized extreme events.
@@ 3.3 Extreme Diagnostics
The preceding analysis has evaluated ensemble weather forecasts from SFNO-BVMC on overall weather. This is necessary but as yet insufficient validation for our main scientific interest in LLHIs. Since extreme weather events are rare in space and time, they contribute relatively little to these scores. Hereafter, we focus on diagnostics specifically designed to validate the performance of SFNO-BVMC on extreme weather.
@@ 3.3.1 Extreme Forecast Index
We generate a model climatology of SFNO-BVMC using the same parameters as ECMWF’s M-Climate, except the SFNO-BVMC M-Climate uses 12 ensemble members, not 11. This is due to the use of centered (positive and negative) bred vector perturbations, which requires an even number of ensemble members. After creating the climatology of SFNO-BVMC, we calculate the CDF of the model climate for each lead time for each grid cell. We use these CDFs to calculate the EFI for the SFNO-BVMC forecasts initialized on each day of summer 2023. Figure 13 visualizes a sample EFI from SFNO-BVMC and IFS four days into a forecast on an arbitrary summer day. The IFS EFI values are directly downloaded from the ECMWF MARS data server. The SFNO-BVMC and IFS EFI values have excellent agreement across the globe (Figure 13). Notable features include pronounced heatwaves over much of Africa, South America, and the Midwest of the United States. The strong El Niño pattern in the tropical Pacific appears in the EFI for both SFNO-BVMC and IFS ENS.
Visually, SFNO-BVMC has a smoother EFI than IFS ENS. This is a consequence of the blurriness of the SFNO 2m temperature predictions. While the embedding dimension and scale factors mitigate this blurriness as much as possible, the SFNO-BVMC model climatology and ensemble forecasts have this artifact. Therefore, the EFI values also appear blurry. Despite this, however, the SFNO EFI can still predict large-scale extremes, and the two models have similar scores on the extreme diagnostics below.
@@
The EFI itself does not measure the accuracy of a forecast; it only measures how extreme or unusual a forecast is by comparing a given forecast to the model climatology. To evaluate the accuracy of the extreme forecast, the EFI is compared to an observational dataset to assess if the extreme forecasts match observations. We follow ECMWF’s validation strategy of using a Receiver Operating Characteristic curve to assess how well the EFI predicts the verification values. The ROC curve could be calculated to assess how well the EFI predicts extreme temperatures.
@@ 3.3.2 Reliability and Discrimination
Reliability diagrams and ROC curves are calculated by comparing two quantities: a binarized ground truth value (1 or 0, for extreme and not extreme) and a continuous ensemble forecast between 1 and 0.
@@
A key validation criterion is the threshold defining extreme vs. not extreme. To enable future comparison with GenCast, we calculate our threshold for extreme temperature using the same definition as Price et al. (2023). Using the years 1992-2016 of ERA5, we calculate the climatological 95th percentile 2m temperature for each grid cell. These percentiles are calculated for each time of day (00:00, 06:00, 12:00, and 18:00 UTC) for each month. This results in 48 different thresholds in total. This definition of extreme accounts for the diurnal and seasonal cycles: an event is considered extreme if it is hot for the time of day and time of year. It also includes warm nighttime temperatures, which has important implications for fire (Balch et al., 2022) and human health (Murage et al., 2017; He et al., 2022), and warm winters, which has important implications for agriculture (Lu et al., 2022). This is a different rationale than defining extreme weather using an absolute temperature threshold or a threshold based only on the summer daily maximum.
@@
Figure 15c shows that SFNO-BVMC and IFS ENS are similarly reliable in their prediction of extreme 2m temperatures at these lead times. To create the reliability diagram in Figure 15a, the ground truth dataset is binarized using the extreme temperature threshold defined described above: one threshold for each month for each time of day. The ensemble prediction is a continuous value indicating the proportion of the ensemble that predicts extreme. This performance is aggregated over all grid cells and all initial times of summer 2023. This results in a reliability diagram, calculated for lead times of 120 hours and 240 hours. We visualize the reliability diagrams at other lead times in the Supplemental Figure 17.
@@
Figure 15b shows that SFNO-BVMC and IFS ENS have a comparable ability to discriminate between extremes and non-extremes. Both ensembles have similar ROC Area Under Curve (AUC) scores, which measure the discrimination of an ensemble. The ROC curve varies the threshold for classifying an event as "extreme" or "not extreme" from 0 to 1: for each threshold, the resulting true positive and false positive rates are plotted. A successful ROC curve would have a 0 false positive rate and 1 true positive rate: the area under such a curve would be 1. To calculate the ROC AUC scores in Figure 15b, we use the EFI. To actually compare the EFI to observations, EFI ROC curves serve as ECMWF’s Supplemental Score on Extremes in their IFS validation (Haiden et al., 2023). The IFS EFI is defined on a daily mean temperature, not a six-hourly temperature. Therefore, in the EFI ROC AUC score in Figure 15b uses a threshold based on daily means. This results in 12 thresholds for extreme weather (one for each month), instead of 48 thresholds (one for each month for each time of day, as in (Price et al., 2023)). Based on the available data on the ECMWF MARS data server, we can only access IFS EFI values until a lead time of 7 days, so we only show IFS scores up to that lead time. At long lead times (approaching 14 days), much of the SFNO-BVMC EFI skill comes from the strong El Niño in summer 2023. Because the EFI is only calculated on data with a daily sampling frequency, Figure 15b necessitated a different extreme threshold than Figures 15a and c. This difference is necessary to enable comparison of EFI ROC curves with Haiden et al. (2023) and extreme diagnostics with Price et al. (2023).
@@ 3.3.3 Threshold-weighted Continuous Ranked Probability Score
We calculate threshold-weighted CRPS (twCRPS) on SFNO-BVMC and IFS ENS as a summary score. Extreme weather events have tremendous societal consequences. Therefore, a natural goal is to validate these weather forecasts specifically on their performance for such extremes. One approach might be to evaluate the forecasts during times of extreme weather. However, Lerch et al. (2017) explain the concept of the forecaster’s dilemma, which is a common pitfall that occurs with this strategy. This dilemma occurs when a forecast is validated on its extreme event forecasts only when those extremes actually happen. With this verification setup, a forecast system can hedge its performance by overpredicting extreme events. Since it is never evaluated during common weather, the forecast would not be penalized for its overly extreme predictions. By construction, statistically proper scoring rules do not allow for such hedging, and twCRPS is one such scoring rule (Gneiting and Ranjan, 2011; Allen et al., 2023).
@@
where $w$ is a weighing function, $X$ is a random variable drawn from the ensemble distribution, $y$ is the verification value, and $v$ is the antiderivative of $w$. We refer the reader to Allen et al. (2022) for further discussion of twCRPS. We choose a weighing function
@@
This weighing function is applied at each grid cell. $t$ is the 95th percentile 2m temperature described above; it is calculated for each time of day for each month.
@@
$\text{twCRPS}(F,y)=\text{CRPS}(F_{t},\text{max}(y,t))$
where $F_{t}$ is the CDF of the transformed ensemble. The transformed ensemble is
@@
for each ensemble member $E_{i}$, where $i$ goes from 1 to $N$ for an ensemble size of $N$. As the CDF of the ensemble, $F_{t}(x)$ is thus calculated as $\frac{1}{N}\sum_{i=1}^{N}1(\text{max}(E_{i},t)\leq x)$. This transformation is described in further detail in Allen et al. (2022).
@@
Figure 15b shows that SFNO-BVMC and IFS have similar twCRPS scores. In fact, SFNO-BVMC outperforms IFS, with a lower twCRPS score. Since this score assesses the prediction of the tails of the distribution, SFNO-BVMC is a trustworthy model for predicting extreme 2m temperature events. The twCRPS has the same units as the standard CRPS; for 2m temperature, the units are degrees Kelvin. However, the values for twCRPS are lower than those for CPRS because the former assigns no penalty for the most common case, when both the ensemble members and the ground truth value are below the threshold. In those cases, the twCRPS score will be 0. Relatedly, the score will be very close to 0 if most (though not all) of the ensemble members predict a non-extreme event and the ground truth is a non-extreme event. This behavior brings down the value of the twCRPS score, compared to the CRPS score.
twCRPS complements other forecast diagnostics, including those specifically focused on extremes. Recently, Ben Bouallègue et al. (2024) validate PanGu weather on extreme weather events, in part by comparing quantile-quantile plots of PanGu, IFS, and ERA5. While these plots compare the aggregate distributions of the forecasts and the truth, they do not assess whether extreme forecasts are collocated (in space and time) with extreme observations. Ben Bouallègue et al. (2024) state that additional diagnostic tools are necessary to evaluate this. twCRPS fills this need, as it focuses on the tails of the ensemble distribution, but it also evaluates whether the forecasts coherently predict extremes at the right space and time.

### v3

@@ 3.2 Spectral Diagnostics
A common issue with deterministic machine learning weather models is that their forecasts tend to be “blurry” (Kochkov et al., 2023). As a metric to measure and quantify this blurriness, existing work compares the spectra of the ML predictions to the spectra of ERA5. The spectral analyses show that ML models have reduced power at small wavelengths compared to ERA5. Deterministic ML models are often trained using the MSE loss function, which strongly penalizes sharp forecasts in the wrong place. This is referred to as the double penalty problem (Mittermaier, 2014), in which an ensemble is penalized once for predicting a storm in the wrong place and another time for missing the correct location of the storm. To avoid the double penalty from the mean squared error, ML models may learn to predict smooth, blurred solutions that appear closer to an ensemble mean (Agrawal et al., 2023; Brenowitz et al., 2024), rather than an individual ensemble member.
Regarding spectral performance, there are two desirable characteristics. These characteristics distinguish an individual ensemble member from an ensemble mean:
  1. 1.
    During the rollout, it is preferable for the spectra of each ensemble member to stay constant with lead time. With this characteristic, each ensemble member maintains a realistic representation of the atmospheric state during the rollout.
  2. 2.
@@
    During the rollout, it is preferable for the spectra of the ensemble mean to realistically degrade with lead time (Bonavita, 2023). As the ensemble members spread more and their trajectories diverge, the ensemble mean should become blurrier. In particular, on synoptic time scales (around 3-5 days), when error growth becomes nonlinear, the IFS ensemble mean displays a sharp decline in power around 1000 kilometer wavelengths (Bonavita, 2023).
On the first characteristic, SFNO-BVMC spectra remain constant through the 360-hour rollout (Figure 10 and Figure 11). This contrasts with GraphCast and AIFS; those deterministic ML models do increasingly blur with lead time (Kochkov et al., 2023; Lang et al., 2024). We hypothesize that SFNO-BVMC spectra remain constant because of our intentional choice not to use autoregressive training.
Through this test, we validate that the individual members’ predictions do not collapse into the ensemble mean. This is a crucial test of the physical fidelity of SFNO-BVMC. Because each SFNO-BVMC ensemble member’s spectrum is constant through the rollout, the ensemble members maintain their ability to resolve extreme weather. If their spectra degraded with lead time, then the forecasts may become too blurry to predict highly localized extreme events. At a lead time of 360 hours, the perturbed members maintain similar spectra as the control member (Figure 11), and at the initial time, they have similar spectral characteristics as the unperturbed ERA5 initial condition (Figure 22).
An important caveat is that even though the spectra are constant during the rollout, they are still somewhat degraded compared to ERA5 (Figure 2b) We have not solved the problem of blurry forecasts entirely. We have minimized it as much as possible by using a large embedding dimension and a small scale factor, which increase the resolution of the latent representation of the input, and by intentionally avoiding multistep finetuning. However, our deterministic training setup still results in blurring with the use of the MSE loss function and large six-hour timesteps, and alleviating this problem is an important avenue for future research.
On the second characteristic, the SFNO-BVMC ensemble mean realistically degrades with lead time: it has a similar ensemble mean spectra as the IFS ensemble mean. Figure 12 shows that the ensemble means of SFNO-BVMC and IFS ENS similarly degrade in power after 24 hours, 120 hours, and 240 hours. For Z500, there is a notable decline in power between lead times of 24 hours and 120 hours. This sharp decline is due to the nonlinear error growth that characterizes forecasts at lead times of 3–5 days. On synoptic scales ($\sim$1000 km in space and 3–5 days in time), SFNO-BVMC’s ensemble mean has a similar decline in power as IFS ENS. This increases our trust that the ensemble members trajectories realistically diverge, and the ensemble is correctly representing synoptic error growth.
These two results pass a crucial test laid out by Bonavita (2023). They originally posed this test comparing the spectra of a deterministic PanGu ML model and the IFS ensemble mean. Despite the blurring in PanGu, they show that a control run of PanGu does not successfully mimic the IFS ensemble mean spectrum. In our work, we have created an ensemble prediction system from multiple deterministic ML models that meets the above two characteristics.
@@ 3.3 Extreme Diagnostics
The preceding analysis has evaluated ensemble weather forecasts from SFNO-BVMC on overall weather. This is necessary but as yet insufficient validation for our main scientific interest in LLHIs. Since extreme weather events are rare in space and time, they contribute relatively little to these scores. Hereafter, we focus on diagnostics specifically designed to validate the performance of SFNO-BVMC on extreme weather. We complement these diagnostics with a case study of the Phoenix 2023 heatwave in Figure 16.
@@ 3.3.1 Extreme Forecast Index
We generate a model climatology of SFNO-BVMC using the same parameters as ECMWF’s M-Climate, except the SFNO-BVMC M-Climate uses 12 ensemble members, not 11. This is due to the use of centered (positive and negative) bred vector perturbations, which requires an even number of ensemble members. After creating the climatology of SFNO-BVMC, we calculate the CDF of the model climate for each lead time for each grid cell. We use these CDFs to calculate the EFI for the SFNO-BVMC forecasts initialized on each day of summer 2023. Figure 13 visualizes a sample EFI from SFNO-BVMC and IFS four days into a forecast on an arbitrary summer day. The IFS EFI values are directly downloaded from the ECMWF data server. The SFNO-BVMC and IFS EFI values have excellent agreement across the globe (Figure 13). Notable features include pronounced heatwaves over much of Africa, South America, and the Midwest of the United States. The strong El Niño pattern in the tropical Pacific appears in the EFI for both SFNO-BVMC and IFS ENS. Visually, SFNO-BVMC has a smoother EFI than IFS ENS. This is a consequence of the blurriness of the SFNO 2m temperature predictions. Despite this, however, the SFNO EFI can still predict large-scale extremes, and the two models have similar scores on the extreme diagnostics below.
@@
The EFI itself does not measure the accuracy of a forecast; it only measures how extreme or unusual a forecast is by comparing a given forecast to the model climatology. To evaluate the accuracy of the extreme forecast, the EFI is compared to an observational dataset to assess if the extreme forecasts match observations. We follow ECMWF’s validation strategy of using a Receiver Operating Characteristic curve to assess how well the EFI predicts the verification values.
@@ 3.3.2 Reliability and Discrimination
Reliability diagrams and ROC curves are calculated by comparing two quantities: a binarized ground truth value (1 or 0, for extreme and not extreme) and a continuous ensemble forecast between 1 and 0. A key validation criterion is the threshold defining extreme vs. not extreme. We calculate our threshold for extreme temperature using the same definition as Price et al. (2023). Using the years 1992-2016 of ERA5, we calculate the climatological 95th percentile 2m temperature for each grid cell. These percentiles are calculated for each time of day (00:00, 06:00, 12:00, and 18:00 UTC) for each month. This results in 48 different thresholds in total. This definition of extreme accounts for the diurnal and seasonal cycles: an event is considered extreme if it is hot for the time of day and time of year. It thus includes warm nighttime temperatures, which have important implications for fire (Balch et al., 2022) and human health (Murage et al., 2017; He et al., 2022), and warm winters, which have important implications for agriculture (Lu et al., 2022). This is a different rationale than defining extreme weather using an absolute temperature threshold or a threshold based only on the summer daily maximum.
@@
Figure 15c shows that SFNO-BVMC and IFS ENS are similarly reliable in their prediction of extreme warm 2m temperatures at lead times of 120 and 240 hours. To create the reliability diagram in Figure 15a, the ground truth dataset is binarized using the extreme temperature threshold defined described above. The "Forecast Probability" is a continuous value from 0 to 1, indicating the proportion of the ensemble that exceeds the threshold. Over all grid cells and initial times of summer 2023, the reliability diagram compares the probabilistic forecasts of extreme events to their actual occurrence. In addition to the lead times in Figure 15c), we visualize the reliability diagrams for other lead times (Supplemental Figure 18) and variables. We show that SFNO-BVMC also performs reliably when forecasting the heat index at lead times of 48, 96, 120, and 240 hours. For 10m wind speed and cold extremes, SFNO-BVMC matches the performance of the IFS ensemble (Figure 19 and Figure 20). However, we also show that at 240 hour lead times, the model is not reliable when it confidently (greater than 50% chance) forecasts wind extremes or cold temperature extremes (see Appendix D and Figure 21 for more discussion). This is an area for future model development.
@@
Next, we assess the ensemble’s discrimination. Figure 15b shows that SFNO-BVMC and IFS ENS have a comparable ability to discriminate between extremes and non-extremes. Both ensembles have similar ROC Area Under Curve (AUC) scores, which measure the discrimination of an ensemble. The ROC curve varies the threshold for classifying an event as "extreme" or "not extreme" from 0 to 1: for each threshold, the resulting true positive and false positive rates are plotted. A successful ROC curve would have a 0 false positive rate and 1 true positive rate: the area under such a curve would be 1. To calculate the ROC AUC scores in Figure 15b, we use the EFI. To actually compare the EFI to observations, EFI ROC curves serve as ECMWF’s Supplemental Score on Extremes in their IFS validation (Haiden et al., 2023). The IFS EFI is defined on a daily mean temperature, not a six-hourly temperature. Therefore, in the EFI ROC AUC score in Figure 15b uses a threshold based on daily means. This results in 12 thresholds for extreme weather (one for each month), instead of 48 thresholds (one for each month for each time of day, as in (Price et al., 2023)). Based on the available data on the ECMWF MARS data server, we can only access IFS EFI values until a lead time of 7 days, so we only show IFS scores up to that lead time. At long lead times (approaching 14 days), much of the SFNO-BVMC EFI skill comes from the strong El Niño in summer 2023. Because the EFI is only calculated on data with a daily sampling frequency, Figure 15b necessitated a different extreme threshold than Figures 15a and c. This difference is necessary to enable comparison of EFI ROC curves with Haiden et al. (2023) and extreme diagnostics with Price et al. (2023).
@@ 3.3.3 Threshold-weighted Continuous Ranked Probability Score
We calculate threshold-weighted CRPS (twCRPS) on SFNO-BVMC and IFS ENS as a summary score. Since extreme weather events have tremendous societal consequences, a natural goal is to validate these weather forecasts specifically on their performance for such extremes. One approach might be to evaluate the forecasts during times of extreme weather. However, Lerch et al. (2017) explain the concept of the forecaster’s dilemma, which is a common pitfall that occurs with this strategy. This dilemma occurs when a forecast is validated on its extreme event forecasts only when those extremes actually happen. With this verification setup, a forecast system can hedge its performance by overpredicting extreme events. Since it would never be evaluated during common weather, the forecast would not be penalized for its overly extreme predictions. By construction, statistically proper scoring rules do not allow for such hedging, and twCRPS is one such scoring rule (Gneiting and Ranjan, 2011; Allen et al., 2023).
@@
where $w$ is a weighing function, $X$ is a random variable drawn from the ensemble distribution, $y$ is the verification value, and $v$ is the antiderivative of $w$. We refer the reader to Allen et al. (2022) for further discussion of twCRPS and its derivation. We choose a weighing function
@@
This weighing function is applied at each grid cell. $t$ is the 95th percentile 2m temperature described above, calculated for each time of day for each month.
@@
$\text{twCRPS}(F,y)=\text{CRPS}(F_{t},\text{max}(y,t))$ (5)
where $F_{t}$ is the CDF of the transformed ensemble.
For each ensemble member $E_{i}$, where $i$ goes from 1 to $N$ for an ensemble size of $N$, the transformed ensemble member is
@@
The CDF of the transformed ensemble, $F_{t}(x)$, is thus calculated as $\frac{1}{N}\sum_{i=1}^{N}1(\text{max}(E_{i},t)\leq x)$, (Allen et al., 2022).
@@
Figure 15b shows that SFNO-BVMC and IFS have similar twCRPS scores, with SFNO-BVMC performing slightly better on this metric. Since this score assesses the prediction of the tails of the distribution, SFNO-BVMC is a trustworthy model for predicting extreme 2m temperature events. The twCRPS has the same units as the standard CRPS; for 2m temperature, the units are degrees Kelvin. However, the values for twCRPS are lower than those for CPRS because the former assigns no penalty for the most common case, when both the ensemble members and the ground truth value are below the threshold. In those cases, the twCRPS score will be 0. Because the ensemble members and the verification are transformed as in Equations 5 and 6, the twCRPS has a lower value than the CRPS score.
twCRPS complements other forecast diagnostics, including those specifically focused on extremes. Recently, Ben Bouallègue et al. (2024) validate PanGu weather on extreme weather events, in part by comparing quantile-quantile plots of PanGu, IFS, and ERA5. While these plots compare the aggregate distributions of the forecasts and the truth, they do not assess whether extreme forecasts are collocated (in space and time) with extreme observations. Ben Bouallègue et al. (2024) state that additional diagnostic tools are necessary to evaluate this. We suggest that twCRPS fills this need, as it focuses on the tails of the ensemble distribution, but it also evaluates whether the forecasts coherently predict extremes at the right space and time.